. A normally distributed population has mean 57,800 and standard deviation 750.

a. Find the probability that a single randomly selected element X of the population is between 57,000 and 58,000.

b. Find the mean and standard deviation of 𝑥̅for samples of size 100.

c. Find the probability that the mean of a sample of size 100 drawn from this population is between 57,000 and 58,000. 


1
Expert's answer
2021-05-07T08:52:26-0400

We are given that a normally distributed population has mean 57,800 and standard deviation 75, i.e.; "\mu = 57800"  and  "\sigma=750" .

Let X = randomly selected element of the population

The z probability is given by;

      Z = "\frac{X-\mu}{\sigma}"  ~ N(0,1) 

a. So, P(57,000 <= X <= 58,000) = P(X <= 58,000) - P(X < 57,000)

P(X <= 58,000) = P(  "\frac{X-\mu}{\sigma}" <= "\frac{58000-57800}{750}"  ) = P(Z <= 0.27) = 0.60642

P(X < 57000) = P(  "\frac{X-\mu}{\sigma}" < "\frac{57000-57800}{750}"  ) = P(Z < -1.07) = 1 - P(Z <= 1.07) = 1 - 0.85769 = 0.14231

Therefore, P(31 < X < 40) = 0.60642 - 0.14231 = 0.46411 .


b. Now, we are given sample of size, n = 100

So, Mean of X, "\overline{X}" = 57,800 same as before

But standard deviation of X, s = "\frac{\sigma}{\sqrt{n}}""\frac{750}{\sqrt{100}}" = 75


c. The z probability is given by;

      Z = "\frac{\overline{X}-\mu}{s}"  ~ N(0,1) 

Now, probability that the mean of a sample of size 100 drawn from this population is between 57,000 and 58,000 = P(57,000 < "\overline{X}" < 58,000)

P(57,000 <= "\overline{X}" <= 58,000) = P("\overline{X}" <= 58,000) - P("\overline{X}" < 57,000)

P("\overline{X}" <= 58,000) = P(  "\frac{\overline{X}-\mu}{s}"  <= "\frac{58000-57800}{75}" ) = P(Z <= 2.67) = 0.99621

P(X < 57000) = P( "\frac{\overline{X}-\mu}{s}" < "\frac{57000-57800}{75}") = P(Z < -10.67) = P(Z > 10.67)

This probability is that much small that it is very close to 0

Therefore, P(57,000 < X bar < 58,000) = 0.99621 - 0 = 0.99621 .